suppose a random sample of 1750 households with television showed that 69.5 percent of them have cable television. Is this sufficient evidence, at the 0.05 significance level, to claim more than two-thirds of all television households have cable televisions?
Answer by stanbon(75887)

suppose a random sample of 1750 households with television showed that 69.5 percent of them have cable television. Is this sufficient evidence, at the 0.05 significance level, to claim more than two-thirds of all television households have cable televisions?
-------------------------------
Ho: p <= 2/3
Ha: p > 2/3 (claim)
---------------------------------
p-hat = 0.695
----
z(0.695) = (0.695-(2/3))/sqrt[(2/3)(1/3)/1750] = 2.5143
-----------------------------
p-value = P(z > 2.5143) = normalcdf(2.5143,100) = 0.006
---------------------------------
Conclusion: Since the p-value is less than 5%, reject Ho.
The test results support the claim.
----------------------------------------